### v2

@@ Abstract
Decentralized Finance (DeFi) has revolutionized lending by replacing intermediaries with algorithm-driven liquidity pools. However, existing platforms like Aave and Compound rely on static interest rate curves and collateral requirements that struggle to adapt to rapid market changes, leading to inefficiencies in utilization and increased risks of liquidations. In this work, we propose a dynamic model of the lending market based on evolving demand and supply curves, alongside an adaptive interest rate controller that responds in real-time to shifting market conditions. Using a Recursive Least Squares algorithm, our controller estimates tracks the external market and achieves stable utilization, while also minimizing risk. We provide theoretical guarantees on the interest rate convergence and utilization stability of our algorithm. We establish bounds on the system’s vulnerability to adversarial manipulation compared to static curves, while quantifying the trade-off between adaptivity and adversarial robustness. Our dynamic curve demand/supply model demonstrates a low best-fit error on Aave data, while our interest rate controller significantly outperforms static curve protocols in maintaining optimal utilization and minimizing liquidations.
Keywords:
Lending Decentralized Finance Recursive Least Squares Adversarial Robustness
1 Introduction
@@
Lending markets in DeFi Decentralized Finance (DeFi) has transformed lending by removing centralized intermediaries like banks, replacing them with transparent, algorithm-driven liquidity pools. Platforms such as Aave [3] and Compound [6] enable lenders to provide capital that borrowers access by pledging collateral. A key goal of these protocols is to maintain stable utilization rates, adjusting interest rates to balance supply and demand [9]. Low utilization results in lower interest rates to encourage borrowing, while rising utilization drives up rates to manage liquidity. Another critical parameter is the collateral factor, which ensures borrower risk is minimized while keeping markets attractive. Setting appropriate collateral levels requires assessing recent price behavior of the collateral asset and the lender’s risk tolerance [12].
@@
Current approaches Current DeFi platforms determine interest rates based on a static function of utilization [8, 7]. This approach uses utilization to represent supply/demand dynamics, market risk, and attractiveness, relying on a manually set, arbitrary interest rate curve. Additionally, collateral factors in these markets are established through a thorough process involving community proposals and reviews [2, 5], which are voted on every few months or so.
@@
Challenges in adapting to market conditions Current DeFi systems are slow to adapt to rapid market changes, leading to potential losses and increased risks due to delayed parameter adjustments, particularly during major price fluctuations. For instance, between August 1 and August 6, 2024, Aave ETH V3 experienced over \$116M in liquidations due to a significant price drop in ETH. Among the affected users, only 45% (\$53M) actively increased their collateral just before being liquidated. It turns out that a 15% increase in the over-collateralization ratio could have nearly halved the number of liquidations for these users (see Figure 1(b)). However, Aave’s slow response in adjusting collateral ratios and liquidation thresholds failed to avert these outcomes. Additionally, Aave has faced challenges in maintaining optimal utilization levels during rapid, unforeseen market changes, particularly when demand spikes due to new yield farming opportunities or similar events (see Figure 1(a)).
@@
Adaptivity versus Robustness Introducing adaptivity into a protocol can increase its vulnerability to manipulation [10], as an adaptive protocol adjusts based on recent market activities, learning from the trades made by borrowers and lenders. However, an adversary can exploit this by distorting the historical trade data, making the protocol respond to false shifts in demand or supply. Therefore, it is crucial to quantify the tradeoff between adaptivity and adversarial robustness before deploying any new lending protocol. Specifically, we assess how much the interest rate can be manipulated when an adversary controls a portion of the total demand and supply.
@@ 1.1 Our Contributions
Agile Interest Rate Controller: (Section 3) We propose an adaptive interest rate controller that adjusts dynamically based on real-time market conditions, ensuring a desired stable utilization in presence of changing environments. The market conditions are estimated in an online manner, using a variant of the Recursive Least Squares algorithm. We provide theoretical guarantees on the controller’s convergence speed to the target interest rate. Additionally, we outline how the algorithm can be augmented with a risk minimization controller, which keeps defaults and liquidations below a desired threshold.
@@ 1.2 Related Work
Collateralized Lending in Traditional Finance Collateralized lending has been extensively studied in traditional finance, where the main challenge is information asymmetry between borrowers and lenders, leading to issues like moral hazard and credit rationing [16, 32]. In contrast, DeFi platforms benefit from shared access to crypto price history, allowing for more transparent, algorithmic risk assessment. The key challenge in DeFi borrow-lending is determining a fair, adaptive interest rate in a rapidly changing, competitive environment, unlike traditional finance where rates change infrequently and competition is limited.
@@
Collateralized Lending in Decentralized Finance and Adaptive Protocols Various models have explored lender and borrower behavior in DeFi, focusing on interest rate equilibria and protocol efficiency. Some models propose equilibrium-based rates but overlook external markets and default risk strategies [20]. Others consider external market dynamics but lack focus on long-term decisions and liquidation risks [30]. Nash equilibrium studies highlight how protocol-driven prices can cause oscillations and require contract adjustments [18]. Empirical data has helped refine models of user behavior [23, 24, 31], and adversarial risks to DeFi protocols have been examined, exposing vulnerabilities [17, 19, 15]. Recent work on adaptive financial mechanisms addresses impermanent loss and arbitrage in market makers [34, 22, 27, 28, 29]. Strategic behavior exploiting static rate curves has also been identified [33]. Protocols like Morpho [11] and Ajna [4] support adaptive rate discovery but require continuous monitoring. Our work is closest to [13] which proposes a two-timescale adaptive lending protocol. However, it uses a simplistic model of user behaviour and only considers shortsighted adversaries. While [13] focuses on a specific demand-supply dynamic yielding a single equilibrium rate, we use a simpler linear demand and supply model with varying parameters, supported by empirical data. This allows for a range of possible equilibrium rates, enabling the protocol to select the one that best achieves the desired utilization.
2 Problem Formulation
2.1 Market Participants and Pool Mechanics
@@
DeFi lending platforms consist of three main participants: borrowers, lenders, and liquidators, who interact via a smart contract, denoted as $\mathcal{P}$  [^1] A more in-depth explanation of pool mechanics is given in Appendix [0.A.1]. These interactions occur in discrete time slots, each corresponding to a block.
Borrowers Borrowers take loans by providing collateral $\mathcal{A}_{c}$, borrowing a different asset $\mathcal{A}_{l}$ supplied by lenders. The loan amount for borrower $i$ at time $t$ is $B_{t}(i)$, with total borrowed amount $B_{t}$. The collateral provided is $C_{t}(i)$ and the total collateral is $C_{t}$. Borrowers must meet the condition $\frac{B_{t}(i)}{C_{t}(i)p_{t}}<c_{t}<1$, where $p_{t}$ is the price of $\mathcal{A}_{c}$ relative to $\mathcal{A}_{l}$ and $c_{t}$ is the protocol’s collateral factor. Interest rate $r_{t}$ is applied to open positions, increasing $B_{t}$ over time.
Lenders Lenders deposit $\mathcal{A}_{l}$ to earn interest from borrowers. The platform gives all accrued interest to lenders without reserves for defaults, meaning lenders directly bear default risk. Lender $i$’s deposit is $L_{t}(i)$ and the total deposit is $L_{t}$. The utilization rate is $U_{t}=\frac{B_{t}}{L_{t}}$, and lenders earn interest $r_{t}$ on the utilized portion. In case of default, a lender’s loss is proportional to their deposit in the pool.
Liquidators A liquidation mechanism activates when the loan-to-value (LTV) ratio exceeds the liquidation threshold $LT_{t}<1$. Liquidators can repay part of the borrower’s debt in $\mathcal{A}_{l}$ and claim a portion of their collateral $\mathcal{A}_{c}$ plus an incentive fee $LI_{t}$, until the LTV ratio falls below $LT_{t}$.
Protocol ($\mathcal{P}$) The smart contract controls the pool parameters $\{r_{t},c_{t},LT_{t},LI_{t}\}$ in each time slot, with an interface for users to interact according to these rules.
2.2 Asset price model
@@
We model the system in discrete time intervals, normalized to the blocktime. The price of the collateral asset $p_{t}$ is tracked from block to block, while the lent asset, $\mathcal{A}_{l}$, is assumed to be a stablecoin with minimal price fluctuation. The collateral asset price $p_{t}$ follows an exogenous geometric Brownian motion with stochastic volatility $\sigma_{t}$, based on the Heston model [25].
$\sigma_{t}^{2}=\sigma_{t-1}^{2}+\kappa(\theta-\sigma_{t-1}^{2})+\xi\sigma_{t-1}\eta_{t},\quad\eta_{t}\sim\mathcal{N}(0,1)$ (1)
$p_{t}=p_{t-1}\exp\left(\left(\mu-\frac{\sigma_{t}^{2}}{2}\right)+\sigma_{t}\varepsilon_{t}\right),\qquad\varepsilon_{t}\sim\mathcal{N}(0,1)$ (2)
@@
2.3 Users’ incentives
Each borrower (lender) $i$ is specified with their time varying demand (supply) capacity denoted by $B_{t}(i)$ ($L_{t}(i)$) and their private value $r^{b}_{t}(i)$ ($r^{l}_{t}(i)$), representing the maximum (minimum) interest rate they are willing to pay (receive from) $\mathcal{P}$ to remain engaged without seeking alternative markets. The rates $r^{b}_{t}(i)$ and $r^{l}_{t}(i)$, referred to as external interest rates in this paper, represent the rates that alternative markets with comparable risk to $\mathcal{P}$ would offer to borrower (or lender) $i$.
@@ 2.3 Users’ incentives
True demand/supply curve The true demand curve is a function of the interest rate, $r$, representing the total demand at that rate if all the borrowers are truthful:
$f(r;\theta_{t})\coloneqq\sum_{i}B_{t}(i)\mathbbm{1}_{(r^{b}_{t}(i)\geq r)}$ (3)
Similarly, the true supply curve represents the total supply at that rate and utilization if all the lenders are truthful:
$g(r\,U;\omega_{t})\coloneqq\sum_{i}L_{t}(i)\mathbbm{1}_{(r^{l}_{t}(i)\leq r\,U)}$ (4)
In this paper, the true demand and supply curves are parameterized by $\theta_{t}$ and $\omega_{t}$, which are subject to temporal variations, reflecting changes in user behavior.
@@
Utility function [^2] Formal details on the utility functions of strategic users are given in Appendix [0.A.2]. Strategic lenders and borrowers decide how much to allocate to the protocol versus external markets based on their expected returns. A strategic lender $i$ will allocate some portion of their supply, $\hat{L}_{t}(i)$, to the protocol while placing the rest in an external market offering a rate $r^{l}_{t}(i)$. The lender’s goal is to maximize their overall returns from both sources over time. Similarly, a strategic borrower $i$ selects how much to borrow, $\hat{B}_{t}(i)$, to minimize their overall borrowing costs. Calculating these optimal decisions can be quite complex, so we simplify the utility functions in certain cases to better analyze user behavior within the protocol.
2.4 Protocol objectives
2.4.1 Demand-Supply balancing
@@
In decentralized finance (DeFi) platforms, maintaining an optimal balance between supply and demand is crucial, and a key metric for balance is the platform’s proximity to the target utilization rate, $U^{*}$ [23, 24]. Low utilization indicates that deposited supply is not being efficiently utilized, while high utilization negatively impacts user experience, as it restricts lenders from withdrawing their funds, effectively locking them up.
@@
To assess the effectiveness of a protocol $\mathcal{P}$ in achieving optimal utilization, we define the optimal interest rate and measure the deviation of the protocol’s interest rate from this target. Specifically, we consider a pool where demand and supply are modeled as noisy versions of $f(r;\theta_{t})$ and $g(rU;\omega_{t})$, respectively with a Gaussian noise with standard deviation $\nu$. At time $t_{0}$, the parameters $\theta_{t}$ and $\omega_{t}$ shift from their initial values to $\theta$ and $\omega$, after which they remain constant. This allows the system to stabilize at a certain utilization $U$ and interest rate $r$. The optimal interest rate $r^{*}_{\text{utl}}$ is defined as:
@@
$r^{*}_{\text{utl}}\coloneqq\arg\min_{r}\left|U-U^{*}\right|\quad\text{subject to}\quad U=\frac{f(r;\theta)}{g(rU;\omega)}$ (5)
@@
We propose a metric called Rate Deviation denoted by $\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*};\{\theta,\omega\})$ to assess the $\mathcal{P}$’s performance in maintaining $U^{*}$:
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*};\{\theta,\omega\})\coloneqq\mathbb{E}\left[\left|r_{t}-r^{*}_{\text{utl}}\right|\right]$ (6)
@@
The expectation is taken over the protocol’s randomness, as well as the noise present in the users’ demand and supply.  [^3] A more pragmatic version of demand-supply balancing that maximizes the revenue of the protocol is formally defined in Appendix [0.A.3] and Appendix [0.A.4]
2.4.2 Adversarial robustness

### v3

@@ Abstract.
Decentralized Finance (DeFi) has revolutionized lending by replacing intermediaries with algorithm-driven liquidity pools. However, existing platforms like Aave and Compound rely on static interest rate curves and collateral requirements that struggle to adapt to rapid market changes, leading to inefficiencies in utilization and increased risks of liquidations. In this work, we propose a dynamic model of the lending market based on evolving demand and supply curves, alongside an adaptive interest rate controller that responds in real-time to shifting market conditions. Using a Recursive Least Squares algorithm, our controller tracks the external market and achieves stable utilization, while also controlling default and liquidation risk. We provide theoretical guarantees on the interest rate convergence and utilization stability of our algorithm. We establish bounds on the system’s vulnerability to adversarial manipulation compared to static curves, while quantifying the trade-off between adaptivity and adversarial robustness. We propose two complementary approaches to mitigating adversarial manipulation: an algorithmic method that detects extreme demand and supply fluctuations and a market-based strategy that enhances elasticity, potentially via interest rate derivative markets. Our dynamic curve demand/supply model demonstrates a low best-fit error on Aave data, while our interest rate controller significantly outperforms static curve protocols in maintaining optimal utilization and minimizing liquidations.
conference: ; ;
1. Introduction
@@
Lending markets in DeFi Decentralized Finance (DeFi) has transformed lending by eliminating centralized intermediaries such as banks, replacing them with transparent, algorithm-driven liquidity pools. Platforms such as Aave (aaveWP, [n. d.]) and Compound (compoundWP, [n. d.]) enable lenders to provide the capital that borrowers can access by pledging collateral. A key goal of these protocols is to maintain stable utilization rates, adjusting interest rates to balance supply and demand (irThumbRule, [n. d.]). Low utilization results in lower interest rates to encourage borrowing, while increased utilization drives up rates to manage liquidity. Another critical parameter is the collateral factor, which ensures that the risk of borrowing is minimized while the markets remain attractive. Setting appropriate collateral levels requires evaluating recent price behavior of the collateral asset and the lender’s risk tolerance (riskColDeFi, [n. d.]).
@@
Current approaches Current DeFi platforms determine interest rates based on a static function of utilization (aaveIR, [n. d.]; compoundIR, [n. d.]). This approach uses utilization to represent supply/demand dynamics, market risk, and attractiveness, relying on a manually set, arbitrary interest rate curve. Additionally, collateral factors in these markets are established through a thorough process involving community proposals and reviews (aaveGov, [n. d.]; compoundGov, [n. d.]), which are voted on every few months or so.
@@
Challenges in adapting to market conditions Current DeFi systems are slow to adapt to rapid market changes, leading to potential losses and increased risks due to delayed parameter adjustments, particularly during major price fluctuations. For instance, between August 1 and August 6, 2024, Aave ETH V3 experienced over \$116M in liquidations due to a significant price drop in ETH. Among the affected users, only 45% (\$53M) actively increased their collateral just before being liquidated. It turns out that a 15% increase in the over-collateralization ratio could have nearly halved the number of liquidations for these users (see Figure 1(b)). However, Aave’s slow response in adjusting collateral ratios and liquidation thresholds failed to avert these outcomes. Additionally, Aave has faced challenges in maintaining optimal utilization levels during rapid, unforeseen market changes, particularly when demand spikes due to new yield farming opportunities or similar events. For example, Figure 1(a) illustrates the utilization of the Aave LUSD-ETH pool during a period of significant volatility in LUSD demand, driven by unstable yield farming opportunities. Although the protocol aims to maintain a target utilization of 0.8, it frequently fails to achieve this target, with deviations persisting for extended periods.
@@
Adaptivity versus robustness Introducing adaptivity into a protocol can increase its vulnerability to manipulation (msGauntletLending23, [n. d.]), as an adaptive protocol adjusts based on recent market activities, learning from the trades made by borrowers and lenders. However, an adversary can exploit this by distorting the historical trade data, making the protocol respond to false shifts in demand or supply. Therefore, it is crucial to quantify the tradeoff between adaptivity and adversarial robustness before deploying any new lending protocol. Specifically, we assess how much the interest rate can be manipulated when an adversary controls a portion of the total demand and supply. We suggest two solutions each operating in a different scale, to mitigate the adversarial manipulations.
@@ 1.1. Our Contributions
Agile interest rate controller: (Section 4.1) We propose an adaptive interest rate controller that adjusts dynamically based on real-time market conditions, ensuring a desired stable utilization in presence of changing environments. The market conditions are estimated in an online manner, using a variant of the Recursive Least Squares algorithm. We provide theoretical guarantees on the controller’s convergence speed to the target interest rate. Additionally, we outline how the algorithm can be augmented with a risk minimization controller, which keeps defaults and liquidations below a desired threshold.
@@
Enhancing the adversarial robustness: (Section 5.3) We propose two complementary methods to mitigate adversarial manipulation, each addressing the issue from a different perspective. The first is an algorithmic approach based on robust recursive least squares, designed to detect and filter out improbable demand and supply patterns. The second focuses on structural market adjustments that encourage greater user elasticity, strengthening the lending pool’s resilience against adversarial manipulation.
@@ 2. Related work
Collateralized lending in traditional finance Collateralized lending has been extensively studied in traditional finance, where the main challenge is information asymmetry between borrowers and lenders, leading to issues like moral hazard and credit rationing (Chan and Kanatas, 1985; Stiglitz and Weiss, 1981). In contrast, DeFi platforms benefit from shared access to crypto price history, allowing for more transparent, algorithmic risk assessment. The key challenge in DeFi borrow-lending is determining a fair, adaptive interest rate in a rapidly changing, competitive environment, unlike traditional finance where rates change infrequently and competition is limited.
@@
Collateralized lending in decentralized finance and adaptive protocols Various models have explored lender and borrower behavior in DeFi, focusing on interest rate equilibria and protocol efficiency. Some models propose equilibrium-based rates but overlook external markets and default risk strategies (Cohen et al., 2023b). Others consider external market dynamics but lack focus on long-term decisions and liquidation risks (Rivera et al., 2023). Nash equilibrium studies highlight how protocol-driven prices can cause oscillations and require contract adjustments (Chiu et al., 2022). Empirical data has helped refine models of user behavior (Group, 2025a, b; Saengchote, 2023), and adversarial risks to DeFi protocols have been examined, exposing vulnerabilities (Chitra et al., 2023; Cohen et al., 2023a; Carre and Gabriel, 2023). Recent work on adaptive financial mechanisms addresses impermanent loss and arbitrage in market makers (Zheng et al., 2021; Goyal et al., 2023; Milionis et al., 2023; Nadkarni et al., 2023, 2024). Strategic behavior exploiting static rate curves has also been identified (Yaish et al., 2023). Protocols like Morpho (morpho, [n. d.]) and Ajna (ajna, [n. d.]) support adaptive rate discovery but require continuous monitoring. Our work is closest to (Bastankhah et al., 2024) which proposes a two-timescale adaptive lending protocol. However, it uses a simplistic model of user behaviour and only considers shortsighted adversaries. While (Bastankhah et al., 2024) focuses on a specific demand-supply dynamic yielding a single equilibrium rate, we use a simpler linear demand and supply model with varying parameters, supported by empirical data. This allows for a range of possible equilibrium rates, enabling the protocol to select the one that best achieves the desired utilization.
3. Problem formulation
3.1. Market participants and pool mechanics
@@
DeFi lending platforms consist of three main participants: borrowers, lenders, and liquidators, who interact via a smart contract, denoted as $\mathcal{P}$. These interactions occur in discrete time slots, each corresponding to a block.
Lenders Lenders deposit asset $\mathcal{A}_{l}$ to earn interest on the amounts borrowed by the borrowers. In this analysis, we focus on a lending protocol where all accrued interest goes directly to the lenders. Unlike many existing platforms, which reserve a portion of the interest for a fund to cover potential defaults and distribute some as rewards to token holders (aaveWP, [n. d.]; compoundWP, [n. d.]), the platform we study has no such reserves. As a result, any borrower default directly reduces the deposits of lenders, transferring the risk of default from the protocol to the lenders. Let $L_{t}(i)$ be the deposit of lender $i$ in $\mathcal{A}_{l}$, and let $L_{t}$ be the total deposits. The utilization rate, $U_{t}=\frac{B_{t}}{L_{t}}$, expresses the fraction of deposits currently borrowed. Lenders earn interest at rate $r_{t}$ on the portion of their deposits that are utilized. In the event of a default, each lender’s loss is proportional to their original deposit relative to the total pool.
3.2. Asset price model
@@
We model the system in discrete time intervals, normalized to the block time. The price of the collateral asset $p_{t}$ is tracked from block to block, while the lent asset, $\mathcal{A}_{l}$, is assumed to be a stablecoin with minimal price fluctuation. The collateral asset price $p_{t}$ follows an exogenous geometric Brownian motion with stochastic volatility $\sigma_{t}$, based on the Heston model (Heston, 1993).
(1) $\sigma_{t}^{2}=\sigma_{t-1}^{2}+\kappa(\theta-\sigma_{t-1}^{2})+\xi\sigma_{t-1}\eta_{t},\quad\eta_{t}\sim\mathcal{N}(0,1)$
(2) $p_{t}=p_{t-1}\exp\left(\left(\mu-\frac{\sigma_{t}^{2}}{2}\right)+\sigma_{t}\varepsilon_{t}\right),\qquad\varepsilon_{t}\sim\mathcal{N}(0,1)$
@@
3.3. Users’ incentives
Each borrower (lender) $i$ is specified with their time varying demand (supply) capacity denoted by $B_{t}(i)$ ($L_{t}(i)$) and their private value $r^{b}_{t}(i)$ ($r^{l}_{t}(i)$), representing the maximum (minimum) interest rate they are willing to pay (receive from) $\mathcal{P}$ to remain engaged without seeking alternative markets. The rates $r^{b}_{t}(i)$ and $r^{l}_{t}(i)$, referred to as external interest rates in this paper, represent the rates that alternative markets with comparable risk to $\mathcal{P}$ would offer to borrower (or lender) $i$. We represent the vector consisting of all borrowers (lenders) external rate by $\mathbf{\hat{r}}^{b}_{t}$ ( $\mathbf{\hat{r}}^{l}_{t}$) and their capacity vector respectively by $\mathbf{\hat{B}}_{t}$ ($\mathbf{\hat{L}}_{t}$)
@@ 3.3. Users’ incentives
True demand/supply curve The true demand curve is the total demand at interest rate $r$ when all the borrowers are truthful, this function is parameterized by the external interest rate and capacity of the borrowers collectively denoted by $\mathcal{B}_{t}\coloneqq\{\mathbf{\hat{B}}_{t},\mathbf{\hat{r}}^{b}_{t}\}$ :
(3) $f(r;\mathcal{B}_{t})\coloneqq\sum_{i}B_{t}(i)\mathbbm{1}_{(r^{b}_{t}(i)\geq r)}$
Similarly, the true supply curve represents the total supply at an effective interest rate when all the lenders are truthful; suppliers’ effective rate is the interest rate they receive on the utilized part of the pool i.e., $r_{\text{eff}}\coloneqq r\cdot U$. The supply function is parameterized by the external interest rate and capacity of the lenders collectively denoted by$\mathcal{L}\coloneqq\{\mathbf{\hat{L}}_{t},\mathbf{\hat{r}}^{l}_{t}\}$ :
(4) $g(r_{\text{eff}};\mathcal{L}_{t})\coloneqq\sum_{i}L_{t}(i)\mathbbm{1}_{(r^{l}_{t}(i)\leq r_{\text{eff}})}$
@@
Utility function Strategic lenders and borrowers decide how much to allocate to the protocol versus external markets based on their expected returns. A strategic lender $i$ will allocate some portion of their supply, $\hat{L}_{t}(i)$, to the protocol while placing the rest in an external market offering a rate $r^{l}_{t}(i)$. The lender’s goal is to maximize their overall returns from both sources over time. Similarly, a strategic borrower $i$ selects how much to borrow, $\hat{B}_{t}(i)$, to minimize their overall borrowing costs. Calculating these optimal decisions can be quite complex, so we simplify the utility functions in certain cases to better analyze user behavior within the protocol.
At any time $t$, a strategic lender $i$ allocates a portion (or potentially all) of their supply $\hat{L}_{t}(i)\leq L_{t}(i)$ to $\mathcal{P}$ and allocate the remaining supply, $L_{t}(i)-\hat{L}_{t}(i)$, to an alternative external market that offers a rate $r^{l}_{t}(i)$. The lender’s objective is to maximize cumulative utility over time, derived from expected interest rates from both $\mathcal{P}$ and the alternative market. Formally, the cumulative utility over time for the lender is expressed as:
(5) $\mathbb{E}\left[\sum_{t=0}^{\infty}\left(\hat{L}_{t}(i)\,r_{t}\,U_{t}+\left(L_{t}(i)-\hat{L}_{t}(i)\right)\,r^{l}_{t}(i)\right)\right]$
(6) $\mathbb{E}\left[\sum_{t=0}^{\infty}\left(\hat{B}_{t}(i)\,r_{t}+\left(B_{t}(i)-\hat{B}_{t}(i)\right)\,r^{b}_{t}(i)\right)\right]$
3.4. Protocol objectives
3.4.1. Demand-supply balancing
@@
In decentralized finance (DeFi) platforms, maintaining an optimal balance between supply and demand is crucial, and a key metric for balance is the platform’s proximity to the target utilization rate, $U^{*}$ (Group, 2025a, b). Low utilization indicates that deposited supply is not being efficiently utilized, while high utilization negatively impacts user experience, as it restricts lenders from withdrawing their funds, effectively locking them up.
@@
To assess the effectiveness of a protocol $\mathcal{P}$ in achieving optimal utilization, we define the optimal interest rate and measure the deviation of the protocol’s interest rate from this target. Specifically, we consider a pool where demand and supply are modeled as noisy versions of $f(r;\mathcal{B}_{t})$ and $g(r_{\text{eff}};\mathcal{L}_{t})$, respectively with a Gaussian noise with standard deviation $\nu$. At time $t_{0}$, the pool variables $\mathcal{B}_{t}$ and $\mathcal{L}_{t}$ shift from their initial values to $\mathcal{B}_{0}$ and $\mathcal{L}_{0}$, after which they remain constant. This allows the system to stabilize at a certain utilization $U$ and interest rate $r$. The optimal interest rate $r^{*}_{\text{utl}}$ is defined as:
@@
(7) $r^{*}_{\text{utl}}\coloneqq\arg\min_{r}\left|U-U^{*}\right|\quad\text{subject to}\quad U=\frac{f(r;\mathcal{B}_{0})}{g(r_{\text{eff}};\mathcal{L}_{0})},\;r_{\text{eff}}\coloneqq U\cdot r$
@@
We propose a metric called Rate Deviation denoted by $\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*})$ to assess the $\mathcal{P}$’s performance in maintaining $U^{*}$:
(8) $\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*})\coloneqq\mathbb{E}\left[\left|r_{t}-r^{*}_{\text{utl}}\right|\right]$
@@
The expectation is taken over the protocol’s randomness, as well as the noise present in the users’ demand and supply.  [^1] A more pragmatic version of demand-supply balancing that maximizes the revenue of the protocol is formally defined in Appendix [A.1] and Appendix [A.2]
3.4.2. Adversarial robustness